Find the coordinate vector [x]B of x relative to the given basis 8-{b1 ,b2} 2 2 2 -14 3 4 [xlB (Simplify your answers.)
Answer #1

simplify in linear equation:

-1a_1+2a_2= 2 ...(1)

-4a_1 -3a_2 = -14 ...(2)

solve the equation:

we get a_1=2, a_2 = 2

thus

\left [ x \right ]_B=\begin{bmatrix} 2\\ 2 \end{bmatrix}